Use a graphing utility to graph the function and approximate its zero to three decimal places.

f(x) = 2e^3x/2 – 803

x=
Answer #1

B fow: 22 -803 342 fa) 2 1803-0 3/2 -4015 = 394-li (401.5) 8oof (3.997,0) 600 400/ / 200 -10 10 20 x - 200 :x=3.997 This is t